It's a brilliantly sunny day. I'm 5'2" tall and my shadow is 6'.
The tree across the street's shadow is 18'. How tall is the tree?

The building across the other street is 30' tall . How tall is it's shadow?

Answer by afilatei(2)

Let Tree's height = T
Let Tree's Shadow = Ts = 18'
My 6' shadow produced by my 5'2" height
Tree's 18' shadow is produced by Tree's ((5'2")/6)*18 = (5'2")*3 = 15'6"
Ans for height of Tree is 15.5' since 12" = 1'.
If Building across other street is 30' tall,
And my 5'2" height produced 6' shadow,
then building's 30' will produce (6/(5'2"))* 30 = 180/(5'2")
= 180/(31/6)
= ((180 * 6) /31)